# Context and Geometry Aware Voxel Transformer for Semantic Scene Completion

### v1

@@ Abstract
Vision-based Semantic Scene Completion (SSC) has gained much attention due to its widespread applications in various 3D perception tasks. Existing sparse-to-dense approaches typically employ shared context-independent queries across various input images, which fails to capture distinctions among them as the focal regions of different inputs vary and may result in undirected feature aggregation of cross-attention. Additionally, the absence of depth information may lead to points projected onto the image plane sharing the same 2D position or similar sampling points in the feature map, resulting in depth ambiguity. In this paper, we present a novel context and geometry aware voxel transformer. It utilizes a context aware query generator to initialize context-dependent queries tailored to individual input images, effectively capturing their unique characteristics and aggregating information within the region of interest. Furthermore, it extend deformable cross-attention from 2D to 3D pixel space, enabling the differentiation of points with similar image coordinates based on their depth coordinates. Building upon this module, we introduce a neural network named CGFormer to achieve semantic scene completion. Simultaneously, CGFormer leverages multiple 3D representations (i.e., voxel and TPV) to boost the semantic and geometric representation abilities of the transformed 3D volume from both local and global perspectives. Experimental results demonstrate that CGFormer achieves state-of-the-art performance on the SemanticKITTI and SSCBench-KITTI-360 benchmarks, attaining a mIoU of 16.87 and 20.05, as well as an IoU of 45.99 and 48.07, respectively. Remarkably, CGFormer even outperforms approaches employing temporal images as inputs or much larger image backbone networks. Code for the proposed method is available at https://github.com/pkqbajng/CGFormer.
@@ 1 Introduction
In this paper, we propose a context and geometry aware voxel transformer (CGVT) to lift the 2D features. We observe that context-dependent query tend to aggregate information from the points within the region of interest. A visual comparison of the sampling points between the context-dependent and context-independent queries is presented in Fig. 3. Thus, before the aggregation of cross-attention, we first utilizes a context aware query generator to predict context-aware queries from the input individual images. Additionally, we extend deformable cross-attention from 2D to 3D pixel space, which allows points ending with similar image coordinates to be distinguished by their depth coordinates, as illustrated in Fig. 1(b). Furthermore, we propose a simple yet efficient depth refinement block to enhance the accuracy of estimated depth probability. This involves incorporating a more precise estimated depth map from a pretrained stereo depth estimation network [37], avoiding the heavy computational burden as observed in StereoScene [15].
@@
  • We propose a content and geometry aware voxel transformer (CGVT) to improve the performance of semantic scene completion. This module initializes the queries based on the content of individual input images and extends the deformable cross-attention from 2D to 3D pixel space, thereby improving the performance of feature lifting.
@@ 2.1 Vision-based 3D Perception
Vision-based 3D perception [22, 23, 29, 18, 11, 43, 55, 16, 42] has received extensive attention due to its ease of deployment, cost-effectiveness, and the preservation of intricate visual attributes, emerging as a crucial component in the autonomous driving. Current research efforts focus on constructing unified 3D representations (e.g., BEV, TPV, voxel) from input images. LiftSplat [34] lifts image features by performing outer product between the 2D image features and their estimated depth probability to generate a frustum-shaped pseudo point cloud of contextual features. The pseudo point cloud features are then splatted to predefined 3D anchors through a voxel-pooling operation. Building upon this, BEVDet [11] extends LiftSplat to 3D object detection, while BEVDepth [19] further enhances performance by introducing ground truth supervision for the estimated depth probability. With the advancements in attention mechanisms [3, 54, 51, 31, 43], BEVFormer [22, 49] transforms image features into BEV features using point deformable cross-attention [54]. Additionally, many other methods, such as OFT [35], Petr [26, 27], Inverse Perspective Mapping (IPM) [9], have also been presented to transform 2D image features into a 3D representation.
@@ 3.2 View Transformation
Context-dependent Query Generation. Previous coarse-to-fine approaches [21, 41, 44, 53] typically employ shared context-independent queries for all the inputs. However, these approaches may overlook the differences among different images and aggregate information from irrelevant areas. In contrast, CGFormer first generates context-dependent queries from the image features using a context aware query generator. To elaborate, the context feature $\mathbf{C}\in\mathbb{R}^{H\times W\times C}$ and depth probability $\mathbf{D}\in\mathbb{R}^{H\times W\times D}$ are first derived from the 2D image feature $\mathbf{F}^{\text{2D}}$, then the query generator $f$ takes $\mathbf{C}$ and $\mathbf{D}$ as inputs to generate voxel queries $\mathbf{V}_{\textbf{Q}}\in\mathbb{R}^{X\times Y\times Z}$
@@
where $(X,Y,Z)$ denotes the spatial resolution of the 3D volume. Benefiting from the initialization, the sampling points of the context-dependent queries usually locate within the region of interest. A comparison between the deformable sampling points of two types of queries is displayed in Fig. 3.
Figure 3: Visualization of sampling locations for different categories (left: car, middle: bicycle, right: building). The yellow dot represents the query point, while the red dots indicate the locations of deformable sampling points. The sampling points of the context-dependent query (a) tend to be distributed within the regions of interest, whereas the points of the context-independent query (b) often include irrelevant locations.
@@ 4.1 Quantitative Results
Table 1: Quatitative results on SemanticKITTI [1] test set. represents the reproduced results in  [12, 52]. The best and the second best results are in bold and underlined, respectively. Our CGFormer outperforms temporal stereo-based (Stereo-T) methods or those methods with larger image backbones in terms of IoU and mIoU.
Method Input Image Backbone IoU mIoU
@@
MonoScene [2] Mono EfficientNetB7 34.16 11.08 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10
TPVFormer [12] Mono EfficientNetB7 34.25 11.26 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50
SurroundOcc [45] Mono EfficientNetB7 34.72 11.86 56.90 28.30 30.20 6.80 15.20 20.60 1.40 1.60 1.20 4.40 14.90 3.40 19.30 1.40 2.00 0.10 11.30 3.90 2.40
OccFormer [52] Mono EfficientNetB7 34.53 12.32 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70
IAMSSC [46] Mono ResNet50 43.74 12.37 54.00 25.50 24.70 6.90 19.20 21.30 3.80 1.10 0.60 3.90 22.70 5.80 19.40 1.50 2.90 0.50 11.90 5.30 4.10
VoxFormer-S [21] Stereo ResNet50 42.95 12.20 53.90 25.30 21.10 5.60 19.80 20.80 3.50 1.00 0.70 3.70 22.40 7.50 21.30 1.40 2.60 0.20 11.10 5.10 4.90
VoxFormer-T [21] Stereo-T ResNet50 43.21 13.41 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70
’ DepthSSC [50] Stereo ResNet50 44.58 13.11 55.64 27.25 25.72 5.78 20.46 21.94 3.74 1.35 0.98 4.17 23.37 7.64 21.56 1.34 2.79 0.28 12.94 5.87 6.23
Symphonize [13] Stereo MaskDINO 42.19 15.04 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00
HASSC-S [41] Stereo ResNet50 43.40 13.34 54.60 27.70 23.80 6.20 21.10 22.80 4.70 1.60 1.00 3.90 23.80 8.50 23.30 1.60 4.00 0.30 13.10 5.80 5.50
HASSC-T [41] Stereo-T ResNet50 42.87 14.38 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10
StereoScene [15] Stereo EfficientNetB7 43.34 15.36 61.90 31.20 30.70 10.70 24.20 22.80 2.80 3.40 2.40 6.10 23.80 8.40 27.00 2.90 2.20 0.50 16.50 7.00 7.20
H2GFormer-S [44] Stereo ResNet50 44.20 13.72 56.40 28.60 26.50 4.90 22.80 23.40 4.80 0.80 0.90 4.10 24.60 9.10 23.80 1.20 2.50 0.10 13.30 6.40 6.30
H2GFormer-T [44] Stereo-T ResNet50 43.52 14.60 57.90 30.40 30.00 6.90 24.00 23.70 5.20 0.60 1.20 5.00 25.20 10.70 25.80 1.10 0.10 0.00 14.60 7.50 9.30
MonoOcc-S [53] Stereo ResNet50 - 13.80 55.20 27.80 25.10 9.70 21.40 23.20 5.20 2.20 1.50 5.40 24.00 8.70 23.00 1.70 2.00 0.20 13.40 5.80 6.40
MonoOcc-L [53] Stereo InternImage-XL - 15.63 59.10 30.90 27.10 9.80 22.90 23.90 7.20 4.50 2.40 7.70 25.00 9.80 26.10 2.80 4.70 0.60 16.90 7.30 8.40
CGFormer (ours) Stereo EfficientNetB7 44.41 16.63 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30
@@
Method Input Image Backbone IoU mIoU
@@
SSCNet [38] LiDAR - 53.58 16.95 31.95 0.00 0.17 10.29 0.00 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.69 0.67
LMSCNet [36] LiDAR - 47.35 13.65 20.91 0.00 0.00 0.26 0.58 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
@@
MonoScene [2] Mono EfficientNetB7 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
TPVFormer [12] Mono EfficientNetB7 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer [52] Mono EfficientNetB7 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
VoxFormer [21] Stereo ResNet50 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
IAMSSC [46] Mono ResNet50 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
DepthSSC [50] Stereo ResNet50 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.71 5.24 3.51
Symphonies [13] Stereo MaskDINO 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (ours) Stereo EfficientNetB7 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
@@ 4.2 Ablation Study
Table 3: Ablation study of the architectural components on SemanticKITTI [1] validation set. CGVT: context and geometry aware voxel transformer. LGE: local and global encoder. 3D-DCA: 3D deformable cross attention. CAQG: context aware query generator. LB: local voxel-based branch. $\mathcal{T}_{XY}$, $\mathcal{T}_{YZ}$, $\mathcal{T}_{XZ}$: planes of the TPV-based branch. DF: dynamic fusion.
@@ 4.3 Qualitative Results
(a) MonoScene [2] (b) VoxFormer [21] (c) OccFormer [52] (d) CGFormer (e) Ground Truth
@@ 5 Conclusions
Appendix A Appendix / supplemental material
In the appendix, we mainly provide implementation details and more experiment results.
A.1 Datasets and Metrics
Datasets. We evaluate our CGFormer on two datasets: SemanticKITTI [1] and SSC-Bench-KITTI-360 [20]. These datasets are derived from the KITTI Odometry [5] and KITTI-360 [24] Benchmarks, respectively. The evaluation focuses on a specific spatial volume: $51.2m$ in front of the car, $25.6m$ to the left and right sides, and $6.4m$ above the car. Voxelization of this volume results in a set of 3D voxel grids with a resolution of $256\times 256\times 32$, where each voxel measures $0.2m\times 0.2m\times 0.2m$. SemanticKITTI provides RGB images with dimensions of $1226\times 370$ as inputs, encompassing 20 unique semantic classes (19 semantic classes and 1 free class). The dataset includes 10 sequences for training, 1 sequence for validation, and 11 sequences for testing. SSC-Bench-KITTI-360 [20] offers 7 sequences for training, 1 sequence for validation, and 1 sequence for testing. It contains 19 unique semantic classes (18 semantic classes and 1 free class), with input RGB images having a resolution of $1408\times 376$.
Metrics. Following previous methods [2, 21, 12], we report the intersection over union (IoU) and mean IoU (mIoU) metrics for occupied voxel grids and voxel-wise semantic predictions, respectively. The interplay between IoU and mIoU offers a comprehensive perspective on the model’s effectiveness in capturing both geometry and semantic aspects of the scene.
A.2 Implementation Details
Network Structures. Consistent with previous researches [12, 2, 45], we utilize a 2D UNet based on a pretrained EfficientNetB7 [39] as the image backbone. The CGVT generates a 3D feature volume with dimensions of $128\times 128\times 16$ and $128$ channels. The numbers of deformable attention layers for cross-attention and self-attention are $3$ and $2$ respectively. We use 8 sampling points around each reference point for the cross and self-attention head. The voxel-based branch of the LGE comprises $3$ stages with 2 residual blocks [8] each. SwinT [28] is employed as the 2D backbone in the TPV-based branch. Both are followed by feature pyramid networks (FPNs) [25] to aggregate multi-scale features for dynamic fusion. The final prediction has dimensions of $128\times 128\times 16$ and is upsampled to $256\times 256\times 32$ through trilinear interpolation to align the resolution with the ground truth.
Training Setup. We train CGFormer for 25 epochs on 4 NVIDIA 4090 GPUs, with a batch size of 4. It approximately consumes 19 GB of GPU memory on each GPU during the training phase. We employ the AdamW [30] optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$ and set the maximum learning rate to $3\times 10^{-4}$. The cosine annealing learning rate strategy is adopted for the learning rate decay, where the cosine warmup strategy is applied for the first $5\%$ iterations.
A.3 Additional Quantitative Results
Results on SemanticKITTI validation set. In Table 6, we present a comparison of CGFormer with state-of-the-art results on the SemanticKITTI validation set. CGFormer outperforms all other methods in terms of both IoU and mIoU. Additionally, it ranks either first or second on most of the classes, demonstrating consistent performance across various semantic categories, as indicated in previous tables.
A.4 Computational Cost
In Table 7, we display the training memory and inference time of CGFormer, along with those of the comparison methods. Additionally, the table includes the corresponding IoU and mIoU metrics for comprehensive comparison. As shown in the table, CGFormer achieves the best performance in terms of both IoU and mIoU, with comparable training memory and inference time.
A.5 Additional Qualitative Results
We offer additional visualization results in Fig.5 and Fig.6. These examples are randomly selected from the SemanticKITTI [1] validation set.
Table 6: Quatitative results on SemanticKITTI [1] validation set. represents the reproduced results in  [12, 52, 50]. The best and the second best results are in bold and underlined, respectively.
Method Input Image Backbone IoU mIoU
$\blacksquare$ road (15.30%)
$\blacksquare$ sidewalk (11.13%)
$\blacksquare$ parking (1.12%)
$\blacksquare$ other-grnd. (0.56%)
$\blacksquare$ building (14.1%)
$\blacksquare$ car (3.92%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ bicycle (0.03%)
$\blacksquare$ motorcycle (0.03%)
$\blacksquare$ other-veh. (0.20%)
$\blacksquare$ vegetation (39.3%)
$\blacksquare$ trunk (0.51%)
$\blacksquare$ terrain (9.17%)
$\blacksquare$ person (0.07%)
$\blacksquare$ bicyclist (0.07%)
$\blacksquare$ motorcyclist (0.05%)
$\blacksquare$ fence (3.90%)
$\blacksquare$ pole (0.29%)
$\blacksquare$ traf.-sign (0.08%)
MonoScene [2] Mono EfficientNetB7 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25
TPVFormer [12] Mono EfficientNetB7 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52
OccFormer [52] Mono EfficientNetB7 36.50 13.46 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86
IAMSSC [46] Mono ResNet50 44.29 12.45 54.55 25.85 16.02 0.70 17.38 26.26 8.74 0.60 0.15 5.06 24.63 4.95 30.13 1.32 3.46 0.01 6.86 6.35 3.56
VoxFormer-S [21] Stereo ResNet50 44.02 12.35 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18
VoxFormer-T [21] Stereo-T ResNet50 44.15 13.35 53.57 26.52 19.69 0.42 19.54 26.54 7.26 1.28 0.56 7.81 26.10 6.10 33.06 1.93 1.97 0.00 7.31 9.15 4.94
DepthSSC [50] Stereo ResNet50 45.84 13.28 55.38 27.04 18.76 0.92 19.23 25.94 6.02 0.35 1.16 7.50 26.37 4.52 30.19 2.58 6.32 0.00 8.46 7.42 4.09
Symphonize [13] Stereo MaskDINO 41.92 14.89 56.37 27.58 15.28 0.95 21.64 28.68 20.44 2.54 2.82 13.89 25.72 6.60 30.87 3.52 2.24 0.00 8.40 9.57 5.76
HASSC-S [41] Stereo ResNet50 44.82 13.48 57.05 28.25 15.90 1.04 19.05 27.23 9.91 0.92 0.86 5.61 25.48 6.15 32.94 2.80 4.71 0.00 6.58 7.68 4.05
HASSC-T [41] Stereo-T ResNet50 44.58 14.74 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10
H2GFormer-S [44] Stereo ResNet50 44.57 13.73 56.08 29.12 17.83 0.45 19.74 28.21 10.00 0.50 0.47 7.39 26.25 6.80 34.42 1.54 2.88 0.00 7.24 7.88 4.68
H2GFormer-T [44] Stereo-T ResNet50 44.69 14.29 57.00 29.37 21.74 0.34 20.51 28.21 6.80 0.95 0.91 9.32 27.44 7.80 36.26 1.15 0.10 0.00 7.98 9.88 5.81
CGFormer (ours) Stereo EfficientNetB7 45.99 16.87 65.51 32.31 20.82 0.16 23.52 34.32 19.44 4.61 2.71 7.67 26.93 8.83 39.54 2.38 4.08 0.00 9.20 10.67 7.84
Table 7: Comparison of training memory and inference time with SOTA methods on the and SemanticKITTI test set. These metrics were measured on the NVIDIA 4090 GPU.
method TPVFormer [12] OccFormer [52] VoxFormer [21] Symphonize [13] StereoScene [15] CGFormer
Training Meomry (M) 18564 18080 18725 17757 19000 19330
Inference Time (ms) 207 199 204 216 258 205
IoU 34.25 34.53 42.95 42.19 43.34 44.41
mIoU 11.26 12.20 12.20 15.04 15.36 16.63
A.6 Limitations
While CGFormer exhibits strong performance on benchmarks, but the accuracy on most of the categories (e.g., person, bicyclist, other vehicle) is unsatisfactory. Improving the performance on these instances could be beneficial for the downstream application tasks. Furthermore, there is a need to explore designing depth estimation networks under multi-view scenarios to extend the geometry-aware view transformation to these scenes. Despite these limitations, we are confident that CGFormer will contribute to advancing the field of 3D perception.
(a) MonoScene [2] (b) VoxFormer [21] (c) OccFormer [52] (d) CGFormer (e) Ground Truth
Figure 5: More qualitative comparison results on the SemanticKITTI [1] validation set.
(a) MonoScene [2] (b) VoxFormer [21] (c) OccFormer [52] (d) CGFormer (e) Ground Truth
Figure 6: More qualitative comparison results on the SemanticKITTI [1] validation set.

### v2

@@ Abstract
Vision-based Semantic Scene Completion (SSC) has gained much attention due to its widespread applications in various 3D perception tasks. Existing sparse-to-dense approaches typically employ shared context-independent queries across various input images, which fails to capture distinctions among them as the focal regions of different inputs vary and may result in undirected feature aggregation of cross-attention. Additionally, the absence of depth information may lead to points projected onto the image plane sharing the same 2D position or similar sampling points in the feature map, resulting in depth ambiguity. In this paper, we present a novel context and geometry aware voxel transformer. It utilizes a context aware query generator to initialize context-dependent queries tailored to individual input images, effectively capturing their unique characteristics and aggregating information within the region of interest. Furthermore, it extend deformable cross-attention from 2D to 3D pixel space, enabling the differentiation of points with similar image coordinates based on their depth coordinates. Building upon this module, we introduce a neural network named CGFormer to achieve semantic scene completion. Simultaneously, CGFormer leverages multiple 3D representations (i.e., voxel and TPV) to boost the semantic and geometric representation abilities of the transformed 3D volume from both local and global perspectives. Experimental results demonstrate that CGFormer achieves state-of-the-art performance on the SemanticKITTI and SSCBench-KITTI-360 benchmarks, attaining a mIoU of 16.87 and 20.05, as well as an IoU of 45.99 and 48.07, respectively. Remarkably, CGFormer even outperforms approaches employing temporal images as inputs or much larger image backbone networks.
@@ 1 Introduction
In this paper, we propose a context and geometry aware voxel transformer (CGVT) to lift the 2D features. We observe that context-dependent query tend to aggregate information from the points within the region of interest. A visual comparison of the sampling points between the context-dependent and context-independent queries is presented in Fig. LABEL:fig:sampling_offset. Thus, before the aggregation of cross-attention, we first utilizes a context aware query generator to predict context-aware queries from the input individual images. Additionally, we extend deformable cross-attention from 2D to 3D pixel space, which allows points ending with similar image coordinates to be distinguished by their depth coordinates, as illustrated in Fig. 1(b). Furthermore, we propose a simple yet efficient depth refinement block to enhance the accuracy of estimated depth probability. This involves incorporating a more precise estimated depth map from a pretrained stereo depth estimation network [33], avoiding the heavy computational burden as observed in StereoScene [14].
@@
  • We propose a context and geometry aware voxel transformer (CGVT) to improve the performance of semantic scene completion. This module initializes the queries based on the context of individual input images and extends the deformable cross-attention from 2D to 3D pixel space, thereby improving the performance of feature lifting.
@@ 2.1 Vision-based 3D Perception
Vision-based 3D perception [22, 23, 26, 18, 10, 38, 55, 15, 37, 17, 51, 41, 45, 46, 49] has received extensive attention due to its ease of deployment, cost-effectiveness, and the preservation of intricate visual attributes, emerging as a crucial component in the autonomous driving. Current research efforts focus on constructing unified 3D representations (e.g., BEV, TPV, voxel) from input images. LiftSplat [30] lifts image features by performing outer product between the 2D image features and their estimated depth probability to generate a frustum-shaped pseudo point cloud of contextual features. The pseudo point cloud features are then splatted to predefined 3D anchors through a voxel-pooling operation. Building upon this, BEVDet [10] extends LiftSplat to 3D object detection, while BEVDepth [19] further enhances performance by introducing ground truth supervision for the estimated depth probability. With the advancements in attention mechanisms [3, 54, 50, 27, 38], BEVFormer [22, 47] transforms image features into BEV features using point deformable cross-attention [54]. Additionally, many other methods, such as OFT [31], Petr [24, 25], Inverse Perspective Mapping (IPM) [8], have also been presented to transform 2D image features into a 3D representation.
@@ 3.2 View Transformation
Context-dependent Query Generation. Previous coarse-to-fine approaches [21, 36, 39, 53] typically employ shared context-independent queries for all the inputs. However, these approaches may overlook the differences among different images and aggregate information from irrelevant areas. In contrast, CGFormer first generates context-dependent queries from the image features using a context aware query generator. To elaborate, the extracted $\mathbf{F}^{2D}$ is fed to the context net and the depth net [19] to generate the context feature $\mathbf{C}\in\mathbb{R}^{H\times W\times C}$ and depth probability $\mathbf{D}\in\mathbb{R}^{H\times W\times D}$, respectively. Then the query generator $f$ takes $\mathbf{C}$ and $\mathbf{D}$ as inputs to generate voxel queries $\mathbf{V}_{\textbf{Q}}\in\mathbb{R}^{X\times Y\times Z}$
@@
where $(X,Y,Z)$ denotes the spatial resolution of the 3D volume. Benefiting from the initialization, the sampling points of the context-dependent queries usually locate within the region of interest. A comparison between the deformable sampling points of two types of queries is displayed in Fig. LABEL:fig:sampling_offset.
(a) RGB (b) VoxFormer [21] (c) OccFormer [52] (d) CGFormer (ours) (e) Ground Truth
Figure 3: Visualization of the sampling locations for different small objects. The yellow dot represents the query point, while the red dots indicate the locations of the deformable sampling points. The sampling points of the context-dependent query (a) tend to be distributed within the regions of interest. Beneficial from this, CGFormer achieve better performance than previous methods.
@@ 4.1 Quantitative Results
Table 1: Quatitative results on SemanticKITTI [1] test set. represents the reproduced results in  [11, 52]. The best and the second best results are in bold and underlined, respectively.
Method IoU mIoU
@@
MonoScene [2] 34.16 11.08 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10
TPVFormer [11] 34.25 11.26 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50
SurroundOcc [40] 34.72 11.86 56.90 28.30 30.20 6.80 15.20 20.60 1.40 1.60 1.20 4.40 14.90 3.40 19.30 1.40 2.00 0.10 11.30 3.90 2.40
OccFormer [52] 34.53 12.32 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70
IAMSSC [42] 43.74 12.37 54.00 25.50 24.70 6.90 19.20 21.30 3.80 1.10 0.60 3.90 22.70 5.80 19.40 1.50 2.90 0.50 11.90 5.30 4.10
VoxFormer-S [21] 42.95 12.20 53.90 25.30 21.10 5.60 19.80 20.80 3.50 1.00 0.70 3.70 22.40 7.50 21.30 1.40 2.60 0.20 11.10 5.10 4.90
VoxFormer-T [21] 43.21 13.41 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70
DepthSSC [48] 44.58 13.11 55.64 27.25 25.72 5.78 20.46 21.94 3.74 1.35 0.98 4.17 23.37 7.64 21.56 1.34 2.79 0.28 12.94 5.87 6.23
Symphonize [12] 42.19 15.04 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00
HASSC-S [36] 43.40 13.34 54.60 27.70 23.80 6.20 21.10 22.80 4.70 1.60 1.00 3.90 23.80 8.50 23.30 1.60 4.00 0.30 13.10 5.80 5.50
HASSC-T [36] 42.87 14.38 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10
StereoScene [14] 43.34 15.36 61.90 31.20 30.70 10.70 24.20 22.80 2.80 3.40 2.40 6.10 23.80 8.40 27.00 2.90 2.20 0.50 16.50 7.00 7.20
H2GFormer-S [39] 44.20 13.72 56.40 28.60 26.50 4.90 22.80 23.40 4.80 0.80 0.90 4.10 24.60 9.10 23.80 1.20 2.50 0.10 13.30 6.40 6.30
H2GFormer-T [39] 43.52 14.60 57.90 30.40 30.00 6.90 24.00 23.70 5.20 0.60 1.20 5.00 25.20 10.70 25.80 1.10 0.10 0.00 14.60 7.50 9.30
MonoOcc-S [53] - 13.80 55.20 27.80 25.10 9.70 21.40 23.20 5.20 2.20 1.50 5.40 24.00 8.70 23.00 1.70 2.00 0.20 13.40 5.80 6.40
MonoOcc-L [53] - 15.63 59.10 30.90 27.10 9.80 22.90 23.90 7.20 4.50 2.40 7.70 25.00 9.80 26.10 2.80 4.70 0.60 16.90 7.30 8.40
CGFormer (ours) 44.41 16.63 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30
@@
Method IoU mIoU
@@
SSCNet [34] 53.58 16.95 31.95 0.00 0.17 10.29 0.00 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.69 0.67
LMSCNet [32] 47.35 13.65 20.91 0.00 0.00 0.26 0.58 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
@@
MonoScene [2] 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
TPVFormer [11] 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer [52] 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
VoxFormer [21] 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
IAMSSC [42] 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
DepthSSC [48] 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.71 5.24 3.51
Symphonies [12] 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (ours) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
@@ 4.2 Ablation Study
Table 3: Ablation study of the architectural components on SemanticKITTI [1] validation set. CGVT: context and geometry aware voxel transformer. LGE: local and global encoder. 3D-DCA: 3D deformable cross attention. CAQG: context aware query generator. LB: local voxel-based branch. $\mathcal{T}_{XY}$, $\mathcal{T}_{YZ}$, $\mathcal{T}_{XZ}$: planes of the TPV-based branch. DF: dynamic fusion. There are 32M predefined parameters.
@@ 4.3 Qualitative Results
(a) RGB (b) VoxFormer [21] (c) OccFormer [52] (d) CGFormer (ours) (e) Ground Truth
@@ 5 Conclusions
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant No. 2023YFB3209800, in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F020003, and in part by the National Natural Science Foundation of China under Grant No. 62301484.